OpenStudy (anonymous):

Need help with 2 questions PLEASE

OpenStudy (anonymous):

For questions 1–2, which graph represents the solution to the given system? 1. y = –x + 2 y = 3x – 1 (1 point) 2. –4x + 3y = –12 –2x + 3y = –18 (1 point)

OpenStudy (anonymous):

1- (3/4,5/4) 2- (-3,-8)

OpenStudy (anonymous):

Thanks can you show me a little bit how you did it??

OpenStudy (anonymous):

sorry, i can't :)

OpenStudy (anonymous):

THe first one is not one of the options but the 2nd one is Thanks

OpenStudy (anonymous):

I gave you best response but can you try #1 again it isnt one of the choices @mar555